# Spinal Line Detection for Posture Evaluation through Training-free 3D Human Body Reconstruction with 2D Depth Images

### v1

@@ 3.1 System Overview
The 3D human posture analysis system proposed in this study consists of a multi-stage pipeline that extracts high-quality 3D meshes and spinal skeletal information from depth images. Figure 1. illustrates the workflow of the entire system, encompassing the full process from input data acquisition to final skeletal prediction.
The system first receives depth images from four directions (front, left, right, and back) to generate 3D Point Clouds for each view. In this process, a mask image is applied for background removal, and points are generated according to the unique coordinate systems of each view. Since the Point Clouds generated from multiple views possess different local coordinate systems, a registration process is essential to integrate them into a single global coordinate system. This system adopts a two-step registration strategy. The first step is global registration, which combines the RANSAC (Random Sample Consensus) algorithm with the FPFH (Fast Point Feature Histograms) feature descriptor. The second step utilizes the ICP (Iterative Closest Point) algorithm for fine registration. After initial registration, the Point-to-Plane ICP algorithm is applied to achieve fine registration. The registration process proceeds in the order of side views (left and right) followed by the back view, structured to minimize cumulative errors by aligning in the order of higher view overlap.
@@ 3.4 Fine Registration (Point to Plane ICP)
These findings align with the theoretical characteristics whereby Point-to-Point ICP is sensitive to slight sliding errors in the surface tangential direction due to isotropic distance minimization, while Point-to-Plane ICP utilizes geometric constraints in the surface normal direction to directly suppress residuals and expand the convergence domain. Additionally, the dataset exhibits positional noise for individual points due to the characteristics of the depth sensor; however, normals are estimated through local averaging, making them relatively stable. For these reasons, the Fine Registration algorithm in this process was adopted as Point-to-Plane ICP.
@@ 3.7 Body Shape Analysis according to Spinal Angle
This calculation procedure allows for the consistent mapping of the criteria of “angles, static holding, and symmetry” from ISO 11226 and the principles of “situational weighting (forward bending, lateral tilting, twisting, and load/frequency/coupling)” from RULA/REBA, using only the 2D key points provided by MediaPipe. Each variable can then be combined with the defined angle ranges (normal/caution/risk) to serve as input indicators for diagnosing the risk of cervical and lumbar discs. Table 4 summarizes the normal/caution/risk ranges for spinal angles as defined in this study. This is the result of redefining the risk level ranges of ISO 11226 and RULA/REBA within the spinal coordinate system used in this experiment.
@@ 4.1 Experimental Environment
The programming language used is Python version 3.10.18, and the main libraries include OpenCV for image processing, TorchGeometry, NumPy ($\geq$1.21.0) and SciPy ($\geq$1.7.0) for numerical calculations, Pillow ($\geq$8.3.0) for image input/output processing, Open3D ($\geq$0.15.0) for 3D data processing, and MediaPipe ($\geq$0.8.9) for human pose recognition. The hardware configuration consists of an NVIDIA A100 GPU with 40GB of memory, 83.5GB of system RAM, and an Intel Xeon-based CPU, with approximately 80GB of disk space used for experiments. The input data comprises depth maps with a resolution of $\times512$, sampled at every 4 pixels (step = 4), with a depth scale adjusted to 100. A 3D integrated model was generated by aligning point clouds from the left, right, and rear views based on the frontal view. The registration process consists of two stages: FPFH-based RANSAC global registration (Global Registration) and Point-to-Plane ICP-based fine registration (Local Refinement). The registration parameters for each view direction are shown in Table 5.
@@
Left 5.0 [ 20.0, 10.0, 5.0] 20,000 X O ($\pm$2°)
Right 3.0 [ 10.0, 5.0, 2.5, 1.0] 100,000 O O ($\pm$2°)
Back 5.0 [ 25.0, 12.0, 6.0] 30,000 X O ($\pm$2°)
@@ 4.2 Results
Figure 10 shows the reconstructed spinal centerline (in orange) and shoulder and pelvic reference lines by the proposed system for both male and female domains. In both the frontal and side views, the spinal centerline is naturally connected along the body’s central axis, and the symmetry of the shoulders and pelvis is stably maintained. Notably, in female subjects, the lumbar lordosis appears more pronounced due to body shape characteristics, while in male subjects, the curvature of the thoracic kyphosis is relatively gentle. These visual results demonstrate that the proposed system can reliably reconstruct the major anatomical axes (spine, pelvis, shoulders) of the human body, even with incomplete depth information from a single viewpoint.
@@
Figure 10: Results of proposed method for two domains (male/female): The orange lines represent the spi-nal, pelvic, and shoulder skeletal lines.

### v2

@@ 3.1 System Overview
The 3D human posture analysis system proposed in this study operates on depth maps generated under controlled conditions and consists of a multi-stage pipeline that extracts high-quality 3D meshes and spinal skeletal information from depth images. The depth maps used in this study were generated using data obtained from the SizeKorea database, which is a nationally managed anthropometric survey that collects and pro-vides standardized high-resolution 3D human body models. [34]
The system first receives depth images from four directions (front, left, right, and back) to generate 3D Point Clouds for each view. In this process, a mask image is applied for background removal, and points are generated according to the unique coordinate sys-tems of each view. Since the Point Clouds generated from multiple views possess different local coordinate systems, a registration process is essential to integrate them into a single global coordinate system. This system adopts a two-step registration strategy. The first step is global registration, which combines the RANSAC (Random Sample Consensus) algorithm with the FPFH (Fast Point Feature Histograms) feature descriptor. The second step utilizes the ICP (Iterative Closest Point) algorithm for fine registration. After the global initialization, we incorporate a quality-driven feedback mechanism that quantitatively assesses geometric consistency among the aligned views and automatically re-runs the alignment with adjusted parameters when the quality is insufficient. After initial registra-tion, the Point-to-Plane ICP algorithm is applied to achieve fine registration. The registra-tion process proceeds in the order of side views (left and right) followed by the back view, structured to minimize cumulative errors by aligning in the order of higher view overlap.
@@ 3.3.2 Hierarchical Registration and Supplementary Design Framework
To improve robustness against occasional failures of global initialization under oc-clusion, noise, or challenging body poses, we introduce a quality-driven feedback mecha-nism that evaluates the geometric consistency of the initially aligned point clouds before proceeding to downstream steps.
After initial alignment, we compute an alignment quality score based on the mean nearest-neighbor distance between sampled points across each aligned view pair. Specif-ically, for a pair of aligned point clouds, we measure the mean nearest-neighbor distance and convert it into an overlap quality in the range [0, 1] by normalizing with the expected body size; the final score is obtained by averaging the overlap quality over all view pairs.
If the alignment quality score falls below a threshold (set to 0.4 in our experiments), the pipeline automatically triggers an adaptive refinement loop that re-invokes the align-ment module with modified settings (e.g., increased RANSAC iterations or alternative ICP strategies) and re-evaluates the result using the same metric. This loop is fully automated and implemented as modular functions (assess_alignment_quality and ap-ply_adaptive_refinement), ensuring that only geometrically consistent reconstructions are passed to mesh generation and LOD-ensemble skeleton analysis.
@@ 3.4 Fine Registration (Point to Plane ICP)
These findings align with the theoretical characteristics whereby Point-to-Point ICP is sensitive to slight sliding errors in the surface tangential direction due to isotropic distance minimization, while Point-to-Plane ICP utilizes geometric constraints in the surface normal direction to directly suppress residuals and expand the convergence domain. Additionally, the dataset exhibits positional noise for individual points due to the characteristics of the depth sensor; however, normals are estimated through local averaging, making them relatively stable. For these reasons, the Fine Registration algorithm in this process was adopted as Point-to-Plane ICP. The statistical significance of these differences is further validated using paired statistical tests in Section 4.3.
@@ 3.7 Body Shape Analysis according to Spinal Angle
This calculation procedure allows for the consistent mapping of the criteria of “angles, static holding, and symmetry” from ISO 11226 and the principles of “situational weighting (forward bending, lateral tilting, twisting, and load/frequency/coupling)” from RULA/REBA, using only the 2D key points provided by MediaPipe. Each variable can then be combined with the defined angle ranges (normal/caution/risk) to serve as input indicators for diagnosing the risk of cervical and lumbar discs. Table 4 summarizes the normal/caution/risk ranges for spinal angles as defined in this study. To enable a structured interpretation of the extracted spinal indicators, this study adopts posture angle ranges derived from the principles of ISO 11226 and RULA/REBA. Within the context of depth-based, non-contact posture analysis, these ranges are utilized to de-scribe relative postural tendencies and to facilitate the stratification of posture conditions into normal, caution, and risk categories. Given the inherent measurement variability and quantization uncertainty associated with depth sensors, the use of stratified angle ranges is intended to enhance robustness and interpretability at the posture risk level rather than precise angle estimation.
Accordingly, the criteria summarized in Table 4 provide a consistent reference framework for ergonomic posture screening and longitudinal monitoring, rather than ab-solute clinical decision boundaries. This represents a redefinition of the risk level ranges specified in ISO 11226 and RULA/REBA within the spinal coordinate system adopted in this experiment. Accordingly, based on these posture assessment principles, the following table presents screening-level angle ranges adapted for depth-based spinal posture analy-sis.
@@ 4.1 Experimental Environment
The programming language used is Python version 3.10.18, and the main libraries include OpenCV for image processing, TorchGeometry, NumPy ($\geq$1.21.0) and SciPy ($\geq$1.7.0) for numerical calculations, Pillow ($\geq$8.3.0) for image input/output processing, Open3D ($\geq$0.15.0) for 3D data processing, and MediaPipe ($\geq$0.8.9) for human pose recognition. All experiments were executed on CPU without GPU acceleration. The hardware configu-ration consists of an Intel Core i5-14600 CPU, 31GB of system RAM, and approximately 80GB of disk space used for experiments. The input data comprises depth maps with a resolution of $\times512$, sampled at every 4 pixels (step = 4), with a depth scale adjusted to 100. A 3D integrated model was generated by aligning point clouds from the left, right, and rear views based on the frontal view. The registration process consists of two stages: FPFH-based RANSAC global registration (Global Registration) and Point-to-Plane ICP-based fine registration (Local Refinement). The registration parameters for each view direction are shown in Table 5.
To ensure bilateral consistency and reproducibility, we apply identical registration parameters to the left and right side views by explicitly reflecting the anatomical bilateral symmetry of the human body. Since the left/right side views exhibit comparable geometric complexity and similar overlap/occlusion patterns, using a standardized RANSAC budg-et avoids view-dependent bias in hypothesis sampling and yields more consistent align-ment outcomes. Accordingly, we set the RANSAC iteration count to 50,000 for both left and right views. In contrast, the back view is integrated after the side views, benefiting from the already aligned lateral structure, which effectively reduces the transformation search space. Therefore, 30,000 iterations are used for the back view to maintain efficiency while preserving robust alignment performance. This symmetric parameter design im-proves bilateral consistency and enhances the stability and reproducibility of the over-all pipeline.
@@
Left 4.0 [ 15.0, 8.0, 4.0, 2.0] 50,000 O O ($\pm$2°)
Right 4.0 [ 15.0, 8.0, 4.0, 2.0] 50,000 O O ($\pm$2°)
Back 5.0 [ 25.0, 12.0, 6.0] 30,000 O O ($\pm$2°)
@@ 4.2 Results
Figure 10 shows the reconstructed spinal centerline (in orange) and shoulder and pelvic reference lines by the proposed system for both male and female domains. The bottom row additionally includes subjects with obese body types, where the overall registration remains stable and the extracted reference lines re-main visually coherent. In both the frontal and side views, the spinal centerline is naturally connected along the body’s central axis, and the symmetry of the shoulders and pelvis is stably maintained. Notably, in female subjects, the lumbar lordosis appears more pronounced due to body shape characteristics, while in male subjects, the curvature of the thoracic kyphosis is relatively gentle. These visual results demonstrate that the proposed system can reliably reconstruct the major anatomical axes (spine, pelvis, shoulders) of the human body, even with incomplete depth information from a single viewpoint.
@@
Figure 10: Results of proposed method for two domains (male/female), including subjects with obese body types (last row): The orange lines represent the spinal, pelvic, and shoulder skeletal lines
4.3 Statistical Analysis
To assess whether the performance differences between Point-to-Point ICP and Point-to-Plane ICP are statistically significant, we performed paired statistical tests on the fine registration outcomes. For each multi-view reconstruction sample (N = 393), we measured ICP fitness and RMSE after local refinement under identical global initialization (RANSAC-FPFH) and correspondence update settings. RMSE and fitness follow the defi-nitions described in Section 3.4.
Normality was evaluated using the Shapiro–Wilk test. Both fitness and RMSE dis-tributions significantly deviated from normality (p < 0.001 for all cases); therefore, we adopted a non-parametric paired comparison. For RMSE, 11 paired measurements con-tained non-finite values (infinite RMSE) due to non-convergent registrations; these pairs were excluded, resulting in 376 valid paired samples for RMSE analysis.
The Wilcoxon signed-rank test showed that the fitness difference was not statistically significant (W = 14,727; p = 9.91e-2), although Point-to-Plane ICP achieved a marginally higher mean fitness (0.6610 vs. 0.6586; +0.37%). In contrast, the RMSE difference was sta-tistically significant on the valid paired set (W = 26,774; p = 5.88e-4, p < 0.001), indicating that Point-to-Plane ICP yields a small but consistent reduction in RMSE across samples (+0.75% improvement; lower is better) over the 376 valid pairs.
We additionally report effect sizes based on Cohen’s d computed from paired differ-ences, which were small for both metrics (fitness: d = 0.0138; RMSE: d = 0.0188). Taken to-gether, despite the marginal effect sizes, Point-to-Plane ICP demonstrates a statistically significant and consistent improvement in RMSE (with a slight gain in mean fitness), and therefore we adopt Point-to-Plane ICP as the fine registration step in the proposed pipe-line.
4.4 Processing Time
To address execution-time considerations for clinical deployment, we report the wall-clock processing time of the proposed pipeline.
All timing measurements were obtained on the same workstation described in Sec-tion 4.1 using Python wall-clock timing. In a full batch run over 131 datasets, the total processing time was 98.5 min (5,910 s), corresponding to 45.1 ± 8.4 s per dataset (min: 28.3 s; max: 67.8 s). This yields a throughput of 79.9 datasets/hour, with an overall processing efficiency of 95.2% as reported by our runtime logger. This runtime includes point cloud generation from multi-view depth images, preprocessing, RANSAC-FPFH global align-ment, Point-to-Plane ICP refinement, mesh construction, LOD generation, and LOD-ensemble skeleton extraction.
The current implementation is not designed for real-time interaction (e.g., per-frame feedback) and is intended for offline reconstruction and clinical screening workflows. Given the measured throughput ( 80 datasets/hour), the proposed system supports batch-mode clinical pipelines, while real-time applications would require further optimi-zation and/or GPU-accelerated implementations. Because the pipeline is modular, users can reduce execution time by lowering the number of LODs evaluated or by relaxing glob-al-registration settings (e.g., fewer RANSAC hypotheses) when a faster but less robust mode is acceptable.
4.5 Limitations
As illustrated in the last row of Figure 10, the proposed pipeline can still achieve sta-ble point-cloud registration even for atypical body types such as obesity. However, it is important to note that the spinal curvature angles reported in this study are not derived from direct observation of the vertebrae, but are instead estimated from the reconstructed external body surface obtained from depth measurements. Consequently, for subjects with substantial soft-tissue thickness or atypical body proportions, the clinical accuracy of curvature-related metrics (e.g., cervical/lumbar lordosis and thoracic kyphosis) cannot be guaranteed. In such cases, the reported angles should be interpreted as screening-level in-dicators rather than definitive diagnostic measurements.
Furthermore, dense hair and loose or thick clothing may introduce depth discontinu-ities and irregular surface microstructures, which can negatively affect surface-normal es-timation and correspondence matching, particularly around the head and shoulder re-gions. To mitigate these effects, standardized acquisition conditions are recommended, such as minimizing hair-induced occlusions and using clothing that preserves the body silhouette.
The proposed approach also assumes quasi-static capture with a consistent neutral pose across the four directional views. Large pose variations, self-occlusions, or inter-view motion reduce the overlapping surface area and can lead to non-convergent registration or inconsistent skeletal line extraction. Therefore, the method may be less reliable for com-plex poses or dynamic movements, and stable acquisition conditions with limited motion are preferred.
Finally, robust performance requires sufficiently dense and well-calibrated depth measurements. Increased sensor noise, missing returns, or low spatial resolution can propagate to point-cloud artifacts, degrading both global and local registration and ulti-mately affecting the reliability of the extracted reference lines and angle estimates. Hence, the proposed pipeline is contingent on a minimum level of depth-sensor quality and sta-ble calibration for consistent results.
Additionally, direct quantitative comparisons with deep learning-based methods (e.g., PointNetLK [28], DCP [29]) on standard benchmarks like SMPL were not conducted in thisstudy. This decision stems from fundamental mismatches between our medical depth data—characterized by four-directional fixed-capture depth maps with specific scale and structural distributions—and the large-scale RGB-D/general human scan da-tasets used to train these models. Applying pre-trained learning-based models to our medical data without extensive fine-tuning or domain adaptation could lead to mislead-ing performance compar-isons. **Instead, we prioritized validation of geometric registra-tion accuracy and stability under clinical acquisition constraints, which aligns with the primary goal of devel oping a training-free, interpretable pipeline for healthcare applica-tions.
Regarding global registration, methods like TEASER++, Faster Global Registration, and Go-ICP—while offering strong theoretical guarantees—were not employed due to their computational overhead, which is prohibitive for real-time clinical workflows. These ap-proaches incur significantly higher processing times (often 10-50x slower) and require ex-tensive preprocessing, making them less suitable for point-of-care applications. Our RAN-SAC-FPFH pipeline achieves comparable robustness with sub-second execution on standard hardware, optimally balancing speed and accuracy under medical acquisition con-straints.
@@ 5 Conclusion
In addition, To further extend this study toward clinical research contexts, future work may incorporate specialized medical imaging data such as X-ray or CT scans. Such integration could enable a more detailed investigation of the relationship between external body geometry and internal skeletal structures. In the present study, the primary focus is placed on validating a depth-based geometric reconstruction framework, while clinical validation and diagnostic applications are left for future investigation.
Consistent with prior studies adopting ISO 11226 and RULA/REBA for vision- or depth-based posture analysis, the posture risk categories defined in this work are intended to support structured interpretation of posture tendencies in non-contact settings. Future studies that further align posture risk criteria with clinically grounded standards may enhance the precision and applicability of depth-based posture risk assessment.